Solve by using addition or substitution. If solution isn't unique then state whether the system is dependent or inconsistent.
10x+2y=7
y=-5x+3

In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate recieve?

The length of a rectangle is 2in. more than twice it's width. If the perimeter of the rectangle is 34 in., find the dimension of the rectangle.

Answer by praseenakos@yahoo.com(507)

1. 10x+2y=7.....(1)
y =5x+3....(2) substituting for y in (1)
we get 10x+2(5x+3)=7
10x+10x+6=7 or 20x =7-6 20x=1 orx=1/20
substituting for x in (2) we get y = 5(1/20)+3 =1/4 + 3 (1+12)/4
y=13/4
therefore the values of x& y are x=1/20 and y=13/4
2.